OpenStudy (anonymous):

Solve the system by elimination { x+y-2z=8 {5x-3y+z=-6 {-2x-y+4z=-13

OpenStudy (kl0723):

x+y-2z=8 -2x-y+4z=-13... Cancel out the y's. ((((-x+2z=-5))))... new equation with 2 variables using eq 1 and 2... lets eliminate the y's x+y-2z=8 5x-3y+z=-6... notice that to eliminate the y's we need to multiply eq #1 times 3 so we get: 3x+3y-6z=24 5x-3y+z=-6... cancel y's out and: ((((8x-5z=18))))... equation with two variables # 2

OpenStudy (kl0723):

now we have a system of two equations with two variables: -x+2z=-5 8x-5z=18... can you solve for one of the variables?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

not sure

OpenStudy (kl0723):

if you multiply equation #1 by 8 you can cancel the x's out... give it a try

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

equation one is -x+2z=-5 right

OpenStudy (kl0723):

yes

OpenStudy (anonymous):

-8x+16z=-40?

OpenStudy (kl0723):

right ;)

OpenStudy (kl0723):

now you can cancel -8x with 8x and add up the rest... try it

OpenStudy (anonymous):

wait wouldn't z cancel out as well?

OpenStudy (anonymous):

or would it be negative z instead of minus

OpenStudy (kl0723):

you will get -8x+16z=-40 8x-5z = 18.... x's cancel out.... z's do not cancel out only x's add up and you get 11z=-22 z=-22/11 z=-2

OpenStudy (kl0723):

we know one variable out of three, we still need two more ;)

OpenStudy (anonymous):

ok do I put in z in the original equations now to find the other variables?

OpenStudy (kl0723):

that's right....I would use -x+2z=-5, it looks easier

OpenStudy (kl0723):

plug in and solve for x ;)

OpenStudy (anonymous):

-x+-4z=-5?

OpenStudy (anonymous):

sorry not -4z but -4

OpenStudy (kl0723):

yes... now solve for x :P

OpenStudy (anonymous):

x=1?

OpenStudy (kl0723):

x=1 :)

OpenStudy (kl0723):

we have one more to go

OpenStudy (anonymous):

ok

OpenStudy (kl0723):

now we can go back to the very first equation and plug in the values for z and for x to solve for our y... try it ;)

OpenStudy (kl0723):

x+y-2z=8 use this one since it looks simple ;)

OpenStudy (anonymous):

um 1+y-(-4)=8?

OpenStudy (anonymous):

not sure what y would equal

OpenStudy (kl0723):

first plug in: (1)+y-2(-2)=8 (1)+y+4=8 y=8-4-1 y=3

OpenStudy (kl0723):

now to prove that this is right: plug in all the values for x,y,z in equation #3 that you were given first, if you get -13 then it is right :P

OpenStudy (kl0723):

what do u get? :P